OpenStudy (anonymous):

If an 900.0 g sample of radium-226 decays to 225.0 g of radium-226 remaining in 3,200 years, what is the half-life of radium-226?

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

@JFraser

OpenStudy (aaronq):

Use the first-order integrated rate law: \(\large ln[A]_t=-kt+ln[A]_o\) where \(ln[A]_o\) is the initial concentration of the substance \(k\) is the decay constant \(t\) is time \(ln[A]_t\) is the concentration of substance at time \(t\) then use: \(\large t_{1/2}=\dfrac{ln(2)}{k}\). \(t_{1/2}\) is the half-life

OpenStudy (anonymous):

@JFraser

OpenStudy (anonymous):

@aaronq i got a question about this

OpenStudy (aaronq):

what is it?

OpenStudy (anonymous):

what is k? actually how do i find k?

OpenStudy (aaronq):

k is the decay constant, you find it by plugging in the rest of the variables. when you have any equation, you need to know all but one of the variables. You plug these in and solve for the unknown - which would be k in this case.

OpenStudy (aaronq):

You may be more familiar with the equaiton written like this \(\large [A]_{t}=[A]_0*e^{-kt}\) but it's the same thing, it's just rearranged

OpenStudy (anonymous):

1600?

OpenStudy (aaronq):

what is 1600?